# Wonderful Team: Zero-Shot Physical Task Planning with Visual LLMs

### v5

@@ 5 Experimental Results
In this section, we evaluate the performance of Wonderful Team across a diverse set of tasks that challenge various aspects of robotic reasoning and manipulation. We address key elements of robotics, including multimodal reasoning, contextual decision-making, and complex spatial planning. Our experiments are categorized into three main groups, each designed to tackle specific challenges while contributing to the broader evaluation of the system’s capabilities.
@@ 5.1 Multimodal Reasoning - Simulated VIMABench
To assess our approach’s ability to understand multimodal prompts, reason through abstract concepts, and follow constraints, we tested it on all 17 tasks from VIMABench (Jiang et al., 2023). Unlike traditional robotics benchmarks, VIMABench offers a broad range of objects and task types (see Figure 7), requiring advanced scene understanding, multimodal comprehension, and precise planning for manipulation.
@@ 5.2 Implicit Goal Inference - Real Robots
To evaluate our framework’s reasoning abilities and visual context understanding in real-world settings, we designed a set of Implicit Goal Inference Tasks, each with four variations, to assess the system’s capacity for long-horizon reasoning and context-aware high-level instructions interpretation (see Figure 8).
@@
1) Fruit Placement: The robot is asked to place each fruit in a color-matched area across various setups using the same general prompt. This task challenges the system to infer the desired placement and sometimes also to identify and correct any initially misplaced fruits (see Figure 8(a)).
@@ 5.3 Spatial Planning - Real Robots
To further challenge our system, we introduced tasks that require precise planning and subgoal management. These tasks test the agent’s ability to produce accurate action sequences and handle dependencies carefully. (see Figure 9).
@@
All four tasks require the robot to generate accurate intermediate subgoals, carefully plan and execute actions within spatial contexts.
@@ 5.4 Results and Discussion
As shown in Figure 10(c), Wonderful Team outperforms baselines across all VIMABench tasks. The Grounding Team and multi-agent structure are crucial; removing or replacing them significantly reduces performance. Methods like Trajectory Generator and our ablation with a separate VLM struggle to detect uncommon objects and lack nuanced reasoning for detection and manipulation. Even with perfect localization (as in NLaP), complex long-horizon planning remains challenging without the multi-agent structure, leading to misinterpretations and errors (Appendix F.1). Ablation studies (Appendix D) confirm the importance of each component in Wonderful Team.
Figure 11: Success rates of Wonderful Team and Trajectory Generator on real-world tasks involving ambiguous instruction tasks and spatial planning tasks.
@@
In real robot tasks with more general instructions (e.g., placing fruits based on color), as shown in Figure 11, Wonderful Team achieved a 100% success rate, while Trajectory Generator significantly struggled due to its separation of reasoning and vision. Trajectory Generator relies on an LLM to extract information from the text prompt, which requires explicit instructions. When multiple objects from the same category (e.g., various fruits) were present without specific identifiers, it failed to distinguish between them. Using only “fruit” as the identifier for LangSAM, it could extract the coordinates of all fruits but could not proceed without knowing each fruit’s identity and color. Since the LLM lacks grounding knowledge and only has access to these coordinates, it fails to perform meaningful reasoning, resulting in ineffective planning and ultimately causing the low success rate.
@@ 6.1 Comparison with Training-Based Methods
In recent years, the machine learning community has often seen new LLMs exceed the performance of previous generation fine-tuned models in zero-shot settings, despite the latter’s advantage of task-specific tuning. To explore this trend in the context of visual LLMs and robotics, we compare Wonderful Team with several methods that were at least partially fine-tuned on high-level planning robotics tasks.
In particular, we compare against: 1) VIMA Jiang et al. (2023) and 2) Instruct2Act Huang et al. (2023a). In Table 1, we consistently see that the advantage of fine-tuning loses out to having a more powerful VLLM.
@@
Task Execution Zero-Shot BC Offline Learning Pre-defined API + One-Shot Ex
@@ 6.2 Limitations: Where does Wonderful Team struggle?
Real-Time Adaptation and Error Recovery: Although the Replanning Agent is designed to address failures post-execution, the framework could be improved with real-time dynamic error detection to catch issues immediately. However, reprocessing parts of or the entire task can be computationally expensive and sometimes impractical, requiring careful system design. This limitation is particularly important in navigation tasks or rapidly changing dynmaic environments, where constant replanning can be costly and reduce applicability. Improving the system’s robustness to environmental variations and enhancing real-time error recovery remain key areas for future work.
High-Level Task Planning: We want to stress that we have only considered high level physical task planning in this paper. Robotics is an exceptionally broad area, and we can not make claims about the effectiveness of our methods on problems such as visual navigation or low-level kinematic control. We have also considered 3D manipulation environments exclusively, and do not make any claims about the applicability of these methods to other domains.
@@ Acknowledgments
We would like to thank Professor Tesca Fitzgerald from the Inquisitive Robotics Lab at Yale University and Professor Yen-Ling Kuo from the UVA Learning and Interactive Robotics Lab for their invaluable support and access to equipment that greatly contributed to our real robot experiments.
@@
  • Appendix B: Prompt and Implementation Details B
  • Appendix C: Experimental Details C
  • Appendix D: Ablations: Are All Parts of Wonderful Team Necessary? D
    • Understanding What Each Part of Wonderful Team Does D.1
  • Appendix E: Ablation Studies: VLLMs’ Spatial Reasoning Limitations and Potentials E
    • Evaluating VLLM’s Spatial Understanding E.1
    • Evaluating VLLMs’ Error Recognition and Correction E.2
  • Appendix F: Comparison with Other Methods F
    • Natural Language as Policies F.1
    • PIVOT F.2
    • Language Models as Zero-Shot Trajectory Generators F.3
    • MOKA F.4
@@ Appendix A Things are Moving Extremely Fast
In the Appendix E, we examine the impact of this rapid progress on the grounding team in particular, and show that older VLLMs often struggled to draw bounding boxes with any regularity, suggesting they lacked the fidelity needed for fine-grained robotic control.
Appendix B Prompt and Implementation Details
@@
B.1 Multi-Agent Architecture
@@
B.2 Wonderful Team Pipeline
@@
B.2.1 High-Level Planning
@@ 3. Missing prerequisite steps
B.2.2 Coordinate-Level Target Location Grounding
@@
B.2.3 Low-Level Action Generation
@@
Appendix C Experimental Details
C.1 Evaluation Protocol
@@
C.2 Multimodal Reasoning - Simulated VIMABench
@@ C.2.1 Task Details
C.2.2 Full Experimental Results
@@
C.3 Implicit Goal Inference - Real Robots
C.3.1 Task Details
@@ C.3.2 Robot Setup
C.3.3 Results
Our real robot experiments demonstrated that our framework successfully completed all three tasks 100% of the time. Note that we did not modify any of the prompt or pipeline moving from simulated VIMABench environment to the tasks on the real robot. It was surprising to us how robust the reasoning and planning capabilities of Wonderful Team are. This section provides qualitative results from these experiments, illustrated in Figures 16, 17, and 18. These figures highlight specific aspects of the tasks, illustrating the effectiveness of our framework. It is important to note that these results only reflect the work of the planning team. The role of the grounding team, locating objects and determining their positions, is crucial for the successful execution of these plans.
@@ C.3.3 Results
In the fruit placement task (Figure 16), we present the final execution plan to illustrate the structure of a complete plan. Due to the straightforward nature of the task, this figure does not include the reasoning process. For the superhero companions and fruit price ranking tasks (Figures 17 and 18), we emphasize the reasoning process and omit the block for the complete final plan for the sake of conciseness. The final plans for these tasks are similar in structure to the fruit placement task, essentially combining the substeps in the execution sequence at the bottom of the figures.
@@
C.4 Spatial Planning - Real Robots
C.4.1 Task Details
@@ C.4.2 Robot Setup
Appendix D Ablation Studies: Are All Parts of Wonderful Team Necessary?
@@
Figure 20 shows the workflow of the complete framework of Wonderful Team. We also provide the full prompt and example input and output corresponding to this workflow chart in Appendix D for more concrete details.
@@
Figure 20 shows the workflow of the complete framework of Wonderful Team. Detailed prompts, input examples, and output corresponding to this workflow can be found in Appendix D.
@@
D.1 Understanding What Each Part of Wonderful Team Does

### v6

@@ 5 Experimental Results
In this section, we evaluate the performance of Wonderful Team across a diverse set of tasks that challenge various aspects of robotic reasoning and manipulation. We address key elements of robotics, including multimodal reasoning, contextual decision-making, and complex spatial planning. Our experiments are categorized into three main groups, each designed to tackle specific challenges and contribute to the broader evaluation of the system’s capabilities.
@@ 5.1 Multimodal Reasoning - Simulated VIMABench
To assess our approach’s ability to understand multimodal prompts, reason about abstract concepts, and follow constraints, we tested it on all 17 tasks from VIMABench (Jiang et al., 2023). Unlike traditional robotics benchmarks, VIMABench offers a broad range of objects and task types (see Figure 7), requiring advanced scene understanding, multimodal comprehension, and precise planning for manipulation.
@@ 5.2 Implicit Goal Inference - Real Robots
To evaluate our framework’s reasoning abilities and visual context understanding in real-world settings, we designed a set of Implicit Goal Inference Tasks, each with four variations, to assess the system’s capacity for long-horizon reasoning and context-aware interpretation of high-level instructions (see Figure 8).
@@
1) Fruit Placement: The robot is asked to place each fruit in a color-matched area across various setups using the same general prompt. This task challenges the system to infer the desired placement and identify and correct any initially misplaced fruits (see Figure 8(a)).
@@ 5.3 Spatial Planning - Real Robots
To further challenge our system, we introduced tasks that require precise planning and subgoal management. These tasks test the agent’s ability to produce accurate action sequences and handle dependencies carefully (see Figure 9).
@@
All four tasks require the robot to generate accurate intermediate subgoals and carefully plan and execute actions within spatial contexts.
@@ 5.4 Results and Discussion
As shown in Figure 10(c), Wonderful Team outperforms baselines across all VIMABench tasks. The Grounding Team and multi-agent structure are crucial; removing or replacing them significantly reduces performance. Methods like Trajectory Generator and our ablation with a separate VLM struggle to detect uncommon objects and lack nuanced reasoning for detection and manipulation. Even with perfect localization (as in NLaP), complex long-horizon planning remains challenging without the multi-agent structure, leading to misinterpretations and errors (Appendix G.1). Ablation studies (Appendix E) confirm the importance of each component in Wonderful Team.
Figure 11: Success rates of Wonderful Team (Ours), Trajectory Generator (TG), and MOKA on real-world robotic tasks, categorized into Implicit Goal Inference Tasks (left) and Spatial Planning Tasks (right). WT achieves perfect success in Implicit Goal Inference Tasks by effectively resolving ambiguous instructions, while TG and MOKA struggle, particularly with conceptual inference. In Spatial Planning Tasks, WT outperforms others, leveraging precise function creation and the verification agent to refine execution. However, all methods face challenges with fine-grained depth estimation (e.g., opening the bottle cap).
@@
In real robot tasks with more general instructions (e.g., placing fruits based on color), as shown in Figure 11, Wonderful Team achieved a 100% success rate, while Trajectory Generator significantly struggled due to its separation of reasoning and vision. Trajectory Generator relies on an LLM to extract information from the text prompt, which requires explicit instructions. When multiple objects from the same category (e.g., various fruits) were present without specific identifiers, it failed to distinguish between them. Using only “fruit” as the identifier for LangSAM, the system could extract the coordinates of all fruits but could not proceed without knowing each fruit’s identity and color. Since the LLM lacks grounding knowledge and only has access to these coordinates, it fails to perform meaningful reasoning, resulting in ineffective planning and ultimately causing the low success rate.
@@ 6.1 Comparison with Training-Based Methods
In recent years, the machine learning community has often seen new LLMs exceed the performance of previous-generation fine-tuned models in zero-shot settings, despite the latter’s advantage of task-specific tuning. To explore this trend in the context of visual LLMs and robotics, we compare Wonderful Team with several methods that were at least partially fine-tuned on high-level planning robotics tasks.
In particular, we compare against: 1) VIMA Jiang et al. (2023) and 2) Instruct2Act Huang et al. (2023a). In Table 1, we consistently see that the advantage of fine-tuning is outweighed by having a more powerful VLLM.
@@
Task Execution Zero-Shot BC Offline Learning Pre-defined API + One-Shot
@@
In addition, we also compare with fine-tuned models RT-1 (Brohan et al., 2022) and OpenVLA (Kim et al., 2024) on a selected subset of the tasks. For a detailed discussion and results of these comparisons, see Appendix G.5.
@@ 6.2 Limitations: Where does Wonderful Team struggle?
Real-Time Adaptation and Error Recovery: Although the Replanning Agent is designed to address failures post-execution, the framework could be improved with real-time dynamic error detection to catch issues immediately. However, reprocessing parts of or the entire task can be computationally expensive and sometimes impractical, requiring careful system design. This limitation is particularly important in navigation tasks or rapidly changing dynamic environments, where constant replanning can be costly and reduce applicability. Improving the system’s robustness to environmental variations and enhancing real-time error recovery remain key areas for future work.
High-Level Task Planning: We want to stress that we have only considered high-level physical task planning in this paper. Robotics is an exceptionally broad area, and we cannot make claims about the effectiveness of our methods on problems such as visual navigation or low-level kinematic control. We have also considered 3D manipulation environments exclusively, and do not make any claims about the applicability of these methods to other domains.
@@ Acknowledgments
We would like to thank Professor Tesca Fitzgerald from the Inquisitive Robotics Lab at Yale University and Professor Yen-Ling Kuo from the UVA Learning and Interactive Robotics Lab for their invaluable support and access to equipment that greatly contributed to our real-world robotics experiments.
@@
  • Appendix B: Scope and Applications B
  • Appendix C: Prompt and Implementation Details C
  • Appendix D: Experimental Details D
  • Appendix E: Ablations: Are All Parts of Wonderful Team Necessary? E
    • Understanding What Each Part of Wonderful Team Does E.1
  • Appendix F: Ablation Studies: VLLMs’ Spatial Reasoning Limitations and Potentials F
    • Evaluating VLLM’s Spatial Understanding F.1
    • Evaluating VLLMs’ Error Recognition and Correction F.2
  • Appendix G: Comparison with Other Methods G
    • Natural Language as Policies G.1
    • PIVOT G.2
    • Language Models as Zero-Shot Trajectory Generators G.3
    • MOKA G.4
    • Fine-tuned Models: RT-1 and OpenVLA G.5
@@ Appendix A Things are Moving Extremely Fast
In the Appendix F, we examine the impact of this rapid progress on the grounding team in particular, and show that older VLLMs often struggled to draw bounding boxes with any regularity, suggesting they lacked the fidelity needed for fine-grained robotic control.
Appendix B Scope and Applications
While many of our tasks focus on table-top pick-and-place scenarios, the design and underlying principles of the Wonderful Team framework extend beyond this domain. This section clarifies the scope of our experiments, showcases the framework’s versatility within table-top manipulation, and explores its potential for broader robotic applications.
B.1 Scope of Experiments
Tasks Beyond Pick-and-Place.
In addition to pick-and-place tasks, we evaluated the framework on tasks where success depends on factors beyond accurate object segmentation. These tasks underscore Wonderful Team’s versatility:
  • Push Tasks (e.g., Sweep Without Exceeding): In this VIMABench task, the robot must push objects into a specified target area while avoiding boundary violations. Beyond accurate object localization, this task introduces two distinct additional challenges: (1) reasoning about the correct starting point outside the objects, requiring logical consideration of the opposite side to enable an effective pushing trajectory, and (2) identifying the target, which is an abstract area rather than a concrete object. Wonderful Team achieves a 90% success rate, showcasing its ability to handle spatial reasoning and abstract target identification alongside precise object detection.
  • Spatial Planning Tasks:
    • Drawing a Five-Pointed Star: This task requires planning and executing a complex, fine-grained trajectory with multiple subgoals, going beyond the simpler two-point trajectories typical of pick-and-place operations. The framework achieves an 80% success rate, successfully coordinating intricate movements. Figure 13 shows the planned trajectory sent to the Supervisor Agent for review prior to execution.
    • Opening a Bottle Cap: This task involves fine-grained manipulation, requiring the robot to understand the 3D scene (e.g., bottle height), precisely position the arm without disturbing the bottle, and apply controlled rotational forces. Wonderful Team achieved a 60% success rate, demonstrating its ability to handle tasks requiring both spatial reasoning and constrained movements. Failure modes were primarily caused by inaccurate depth estimation from the depth camera, leading to the bottle being knocked over—a scenario that is particularly challenging to recover from.
Figure 13: Execution of the drawing task: (Left) Prompted task setup; (Center) Plan generated by Wonderful Team; (Right) Successful execution of the task.
Summary of Our Scope.
Our experiments primarily focus on structured table-top environments with third-person view images. Within this controlled setup, the Wonderful Team framework demonstrates effectiveness across a variety of tasks, including pick-and-place, push, trajectory planning, and rotational manipulation. These results highlight the framework’s versatility and its capacity to extend to a broader range of robotic applications.
B.2 Potential Applications Beyond Table-Top Manipulation
Although our research primarily focused on structured table-top scenarios, the modular design of the Wonderful Team framework, built on generalizable models like GPT-4o, demonstrates significant adaptability. With its self-correction mechanism, failure recovery capabilities, and a grounding team to handle spatial reasoning, the framework can be extended to a wide range of tasks with minimal modifications to the prompts and agent components.
Generalization to Diverse Environments.
Without altering the core framework, Wonderful Team can seamlessly adapt to tasks beyond table-top manipulation. Its ability to perform complex reasoning while extracting accurate positional information from diverse images allows it to handle scenarios involving unstructured and noisy environments.
For example, in the task illustrated in Figure 14, the robot is prompted with: “Plan a dish using ingredients in the cabinet.” This task challenges the robot to reason through a cluttered cabinet with diverse ingredients, some partially occluded, such as the soy sauce. Despite these challenges, Wonderful Team successfully generates a plan (e.g., “Prepare miso soup using miso paste, chicken stock, garlic powder, and soy sauce”), identifies the necessary ingredients, and accurately localizes their positions. This demonstrates the framework’s robustness in complex reasoning and precise localization, even in cluttered and occluded settings.
Figure 14: WT results on reasoning and coordinate extraction in an unstructured cabinet task. WT generates a plan based on the given prompt, identifies the required ingredients, and accurately localizes their positions.
A key challenge in extending beyond table-top environments lies in translating 2D coordinates into 3D world actions. While methods like camera pose estimation or depth cameras can be used, many real-world settings lack precise camera pose data, and depth estimation can introduce inaccuracies. These limitations make it difficult to integrate coordinate-based outputs with action APIs. As our focus is on structured table-top scenarios, we did not extensively address these complexities.
Compared to frameworks like MOKA (Fang et al., 2024), which utilize a separate vision segmentation model for object identification, Wonderful Team exhibits significantly improved accuracy and robustness in handling abstract prompts in noisy environments with diverse object appearances and layouts. GroundingDINO struggles with object identification in noisy environments containing diverse objects with text-heavy labels and varied appearances. As illustrated in Figure 15, it misidentifies corn starch as soy sauce, black pepper as garlic powder, and olive oil as shiro miso, with only the chicken stock correctly identified. Such critical mislabeling severely hampers downstream reasoning and planning tasks, highlighting its limitations in handling complex, cluttered environments.
Figure 15: (a) Segmentation Results from GroundingDINO (used in MOKA) using the same target list as in Figure 14. 3 out of 4 items are misidentified. (b) A variation of prompts using simple descriptions for the soy sauce. Notably, even using an exhaustive list of keywords, it still fails, which makes it almost impossible to recover even with VLLM corrections.
With Wonderful Team’s more robust reasoning and localization capabilities compared to models specifically trained for predefined object detection and grounding tasks, it also demonstrates significant potential for data labeling tasks requiring both reasoning and spatial localization. The high-quality annotations generated by Wonderful Team can potentially serve as valuable training data for future supervised models.
B.3 Navigation
The cabinet example (Figure 14) demonstrates WT’s ability to handle first-person perspectives in cluttered settings, akin to inputs used in navigation tasks. Building on this, we explore its capabilities with a constructed top-down map for navigation tasks, assuming such a map is available, with the following adaptations shown in Figure 16:
  • Obstacle Awareness: Emphasized detecting and incorporating obstacles.
  • Path Planning Module: Replaced the original action conversion step with a path planning algorithm to generate collision-free trajectories based on the start location, goal location, and detected obstacles.
  • Prompt: Slightly adjusted prompts to focus on trajectory planning rather than manipulations.
Figure 16: Adapting the Wonderful Team framework for navigation tasks. (a) Original framework for table-top tasks and (b) adapted framework for navigation, incorporating obstacle and a path planning algorithm.
In our toy navigation setup shown in Figure 17, the system is given the prompt
“The home robot initial location is marked by the orange icon. Bring the magazine on the coffee table to the kitchen counter.”
Figure 17: Top-down view of a home robot environment for a navigation task. The orange icon on the top left corner represents the initial robot location.
Figure 18: Visualization of Wonderful Team’s results for a navigation task, showcasing identified target locations, obstacle positions, and the planned collision-free trajectory.
Wonderful Team’s execution results are shown in Figure 18. The framework first generates a high-level plan consisting of sequential subgoals, such as “navigate to (x, y)” or “place (magazine at (x’, y’))”. These subgoals assume predefined action primitives and are further refined through two critical components:
Target Location Identification: Target locations are determined by converting bounding box information into feasible robot target positions. This step is critical, as an incorrect position—such as one inside or to the left side of the kitchen counter’s bounding box—would result in an infeasible trajectory. Accurate reasoning about the spatial relationship between the robot’s starting position and the target object is essential in the reasoning process.
Obstacle Location Identification: Obstacles are identified as bounding boxes that define “DO NOT ENTER” zones. The entire area within each obstacle’s bounding box is avoided during path planning to ensure a collision-free trajectory.
Figure 18 illustrates the framework’s ability to adapt to a structured home environment by generating obstacle-aware paths. While this toy navigation experiment demonstrates the feasibility of adapting Wonderful Team for navigation tasks, it represents a simplified setting and does not encompass the complexities of full navigation problems, which are beyond the scope of this work.
For more complex navigation environments, the Memory Agent could be extended with a hierarchical memory structure, allowing it to store and recall room-based and object-based spatial information. This extension would enable more sophisticated planning in dynamic and unstructured environments.
Appendix C Prompt and Implementation Details
@@
C.1 Multi-Agent Architecture
@@
C.2 Wonderful Team Pipeline
@@
C.2.1 High-Level Planning
@@ 3. Missing prerequisite steps
C.2.2 Coordinate-Level Target Location Grounding
@@
C.2.3 Low-Level Action Generation
@@
Appendix D Experimental Details
D.1 Evaluation Protocol
@@
D.2 Multimodal Reasoning - Simulated VIMABench
@@ D.2.1 Task Details
D.2.2 Full Experimental Results
@@
D.3 Implicit Goal Inference - Real Robots
D.3.1 Task Details
@@ D.3.2 Robot Setup
D.3.3 Results
Our real robot experiments demonstrated that our framework successfully completed all three tasks 100% of the time. Note that we did not modify any of the prompt or pipeline moving from simulated VIMABench environment to the tasks on the real robot. It was surprising to us how robust the reasoning and planning capabilities of Wonderful Team are. This section provides qualitative results from these experiments, illustrated in Figures 22, 23, and 24. These figures highlight specific aspects of the tasks, illustrating the effectiveness of our framework. It is important to note that these results only reflect the work of the planning team. The role of the grounding team, locating objects and determining their positions, is crucial for the successful execution of these plans.
@@ D.3.3 Results
In the fruit placement task (Figure 22), we present the final execution plan to illustrate the structure of a complete plan. Due to the straightforward nature of the task, this figure does not include the reasoning process. For the superhero companions and fruit price ranking tasks (Figures 23 and 24), we emphasize the reasoning process and omit the block for the complete final plan for the sake of conciseness. The final plans for these tasks are similar in structure to the fruit placement task, essentially combining the substeps in the execution sequence at the bottom of the figures.
@@
D.4 Spatial Planning - Real Robots
D.4.1 Task Details
@@ D.4.2 Robot Setup
Appendix E Ablation Studies: Are All Parts of Wonderful Team Necessary?
@@
Figure 26 shows the workflow of the complete framework of Wonderful Team. We also provide the full prompt and example input and output corresponding to this workflow chart in Appendix E for more concrete details.
@@
Figure 26 shows the workflow of the complete framework of Wonderful Team. Detailed prompts, input examples, and output corresponding to this workflow can be found in Appendix E.
@@
E.1 Understanding What Each Part of Wonderful Team Does